OpenStudy (anonymous):

3y=9x+15 find x and y intercepts thank you :)

OpenStudy (anonymous):

y=mx+b Divide both sides by 3 ;-)

OpenStudy (anonymous):

"b" is your y-intercept, when x = 0 that's what 'y' is

OpenStudy (anonymous):

@agentx5 so x is -5/3

OpenStudy (anonymous):

@goformit100 thanks for the medal :)

OpenStudy (lgbasallote):

x-intercept let y = 0 then solve for x when y = 0 0 = 9x+15 -15 = 9x x = -5/3 right

OpenStudy (anonymous):

ok and y is....... ?

OpenStudy (anonymous):

5?

OpenStudy (lgbasallote):

yup

OpenStudy (anonymous):

k omg thanks so the coordinates are -5/3,0 and 0,5 ?